Let X be a random variable with range {0, 1, 2, . . .}. The discrete version of the formula from problem 2.14 in the textbook, i.e. 2.14 (b), can be written

EX = ΣP(X 2 k). k=1

Use this formula to solve the following problem.

A fair dice is thrown n times. The sample space is S = {1, 2, . . . , 6} n , the outcomes are of the form s = (s1, . . . , sn) ∈ S. Let Yn denote the largest of the results thrown, i.e. Yn is a r.v. with Yn(s1, . . . , sn) = max1≤k≤n sk. Find EYn and show limn→∞ E(Yn) = 6.
Answer #1

Following images have the solution

Y = mare Sue п where sue value of die at seth thron. So can tone valves 1, 2, 3, 4, 5, 67 Now fooryo & {1,2,.,6},

KH K= 1 I È(7-6) n in+ 2 tnton Then PCY Zy) = 1-P (Y CY) - - Now, P(Yn (Y). zl(max Suncy). к l (all of Se cy) - P(s, &Y, Szcy

P(, 29 ) = K ( S < 9 ) s, y = s, e(, , , 211] - f (s, <) - - 6 1. Р (Ү7 )) - - - пEC , ) = P (Y, , K) ИС2. E(Yu) = 6 - 2.5 tot (5.77 Now all of 7 , 3 , 25 anec 1 Los 15);- all convenge to zero os n-D ha انه(-۲ + (-)- 6 مه